Find the length of the arc of the curve from point P to point Q. *2 = (y - 3)3, P(1,9), Q(8, 12)
Answer #1

Anss E14-8) 3 Pl1,9) @ 18,12) th of arc is leangth s = ds - + dro dy do au? - 14-813 213 4-8 = x 213 y = 8+ x dy 26 dae 8 S=S = sa du del 2. 6. 3+1 I X x 18 - +1 312 8 213 tox 9x +4 18 312 312 Il 27 / (9x4+4) 19+4) 3 [1024 are length 5 = √ [ 4² - 13thumbs up